OpenStudy (anonymous):

how many moles of carbon monoxide are needed to react completely with 1.75 moles of iron oxide

OpenStudy (anonymous):

Can you write a balanced chemical reaction equation of iron oxide and carbon monoxide?

OpenStudy (unklerhaukus):

is that Iron (III) oxide?

OpenStudy (anonymous):

Most likely. It's the most common form

OpenStudy (unklerhaukus):

Iron (III) oxide\[{Fe_2O_3}_{(s)}+3CO_{(g)} \rightarrow 2Fe_{(s)} +{3CO_2}_{(g)}\] Iron (II) oxide \[FeO_{(s)} +CO_{(g)} \rightarrow Fe_{(s)}+{CO_2}_{(g)}\] Iron (II,III) oxide \[{Fe_3O_4}_{(s)} +4CO_{(g)} \rightarrow 3Fe_{(s)} +4{CO_2}_{(g)}\]

OpenStudy (unklerhaukus):

so your answer would be \[3 \times 1.75\]\[1 \times 1.75\]or\[4 \times 1.75 \]respectively